# MAESTRO: Multi-modal Adaptive Ensemble for Spectro-Temporal Robust Optimization

### v1

@@ Abstract.
Timely and robust influenza incidence forecasting is critical for public health decision-making. To address this, we present MAESTRO, a Multi-modal Adaptive Ensemble for Spectro-Temporal Robust Optimization. MAESTRO achieves robustness by adaptively fusing multi-modal inputs-including surveillance, web search trends, and meteorological data-and leveraging a comprehensive spectro-temporal architecture. The model first decomposes time series into seasonal and trend components. These are then processed through a hybrid feature enhancement pipeline combining Transformer-based encoders, a Mamba state-space model for long-range dependencies, multi-scale temporal convolutions, and a frequency-domain analysis module. A cross-channel attention mechanism further integrates information across the different data modalities. Finally, a temporal projection head performs sequence-to-sequence forecasting, with an optional estimator to quantify prediction uncertainty. Evaluated on over 11 years of Hong Kong influenza data (excluding the COVID-19 period), MAESTRO shows strong competitive performance, demonstrating a superior model fit and relative accuracy, achieving a state-of-the-art R² of 0.956. Extensive ablations confirm the significant contributions of both multi-modal fusion and the spectro-temporal components. Our modular and reproducible pipeline is made publicly available to facilitate deployment and extension to other regions and pathogens.Our publicly available pipeline presents a powerful, unified framework, demonstrating the critical synergy of advanced spectro-temporal modeling and multi-modal data fusion for robust epidemiological forecasting.
@@
copyright: acmlicensedjournalyear: 2025doi: XXXXXXX.XXXXXXXjournal: TISTjournalvolume: 37journalnumber: 4article: 111publicationmonth: 8ccs: Applied computing Health informaticsccs: Computing methodologies Supervised learning by regressionccs: Computing methodologies Neural networksccs: Applied computing Forecastingccs: Mathematics of computing Time series analysis
@@ 1. Introduction
Seasonal influenza exerts a substantial burden on healthcare systems and society. Accurate short-horizon forecasting of influenza incidence supports resource allocation, early warning, and public communication. However, surveillance data alone may be noisy, lagged, or incomplete, while exogenous drivers such as human information-seeking behavior and weather dynamics evolve at differing temporal scales. This motivates multi-modal learning methods that can fuse heterogeneous signals and capture spectro-temporal patterns.
We propose MAESTRO, a Multi-modal Adaptive Ensemble for Spectro-Temporal Robust Optimization. MAESTRO achieves this by adaptively integrating a suite of powerful components: (i) series decomposition into trend and seasonal components; (ii) Transformer encoders for long-range dependencies; (iii) a Mamba state-space block for efficient sequential modeling; (iv) an optional Bi-directional LSTM (BiLSTM) module to capture complementary temporal patterns; (v) multi-scale temporal convolutions for local feature extraction; (vi) frequency-domain learning for periodicity; and (vii) cross-channel attention for inter-variable coupling. The outputs of these components are dynamically fused before a final temporal projection head performs sequence-to-sequence forecasting. An optional uncertainty estimator can also provide calibrated standard deviations for the predictions.
To validate our approach, MAESTRO was rigorously evaluated on a comprehensive real-world dataset from Hong Kong, spanning over a decade of surveillance, Google Trends, and meteorological data (excluding the COVID-19 pandemic period from training to ensure data integrity). Moreover, We conducted extensive ablation studies to systematically verify the contributions of each architectural component and data modality, providing a clear understanding of the model’s inner workings.
Our work introduces a novel, unified spectro-temporal architecture that synergistically couples series decomposition, state-space modeling, attention mechanisms, and frequency-domain learning to specifically address the complexities of epidemiological forecasting. Moreover, we propose an effective multi-modal integration strategy, utilizing cross-channel attention and adaptive fusion to cohesively merge heterogeneous data sources like epidemiological surveillance, web search trends, and weather. What’s more, through a comprehensive empirical study on influenza in Hong Kong, we not only demonstrate the model’s state-of-the-art accuracy and robustness but also ensure full reproducibility by making the entire pipeline publicly available, thereby facilitating its adoption and extension to other diseases and regions.
@@ 2. Related Work
Our work, MAESTRO, addresses this gap. It proposes a novel, modular architecture that synergistically combines time series decomposition, state-space modeling (Mamba), and frequency-domain learning. Crucially, it employs a cross-channel attention mechanism to robustly fuse multi-modal signals, positioning it as a comprehensive and powerful tool for modern epidemiological forecasting.
@@ 3.1. Problem Formulation
Given a multivariate time series $\mathbf{X}\in\mathbb{R}^{T\times D}$ (surveillance, trends, weather) and target $y\in\mathbb{R}$ (influenza positive rate), we predict $y_{t+1:t+H}$ from $\mathbf{X}_{t-L+1:t}$ with horizon $H$ and window $L$.
@@ 3.2. Notation and Assumptions
We denote elementwise (Hadamard) product by $\odot$ and convolution by $\ast$; $\mathcal{F}(\cdot)$ and $\mathcal{F}^{-1}(\cdot)$ are the 1D FFT and inverse FFT along time. Modalities are indexed by $m\in\{1,\dots,M\}$, channels by $i\in\{1,\dots,C\}$, and time by $t\in\{1,\dots,T\}$. Hidden width is $d$. The moving-average operator $\mathcal{D}_{K}$ is a length-$K$ low-pass filter (trend), and $I-\mathcal{D}_{K}$ yields the high-pass residual (seasonal). We use $\operatorname{softmax}$ for normalized exponentiation over the specified axis, $\operatorname{softplus}(x)=\log(1+e^{x})$ for nonnegativity, $\operatorname{pool}$ for temporal pooling (e.g., mean/max), and $\operatorname{MLP}$ for a small feed-forward network. Unless stated, $\|\cdot\|$ denotes the Euclidean/operator norm and $\rho(\cdot)$ the spectral radius.
@@ 3.3. Architecture Overview
    Series Decomposition: If the use_decomp flag is enabled, the input data is first decomposed into trend and seasonal components using a moving average filter provided by the SeriesDecomposition module.
@@
This modular and hierarchical architecture allows MAESTRO to effectively integrate diverse data modalities and capture a wide range of spectro-temporal dynamics. In summary, a moving-average block decomposes inputs into seasonal and trend components. These streams are then embedded and encoded with Transformer encoders. We add residual Mamba state-space, multi-scale temporal convolutions, and a frequency-domain module (learnable filters in FFT space). Cross-channel attention captures inter-variable dependencies. Finally, a linear temporal projection maps encoder outputs to the prediction horizon $H$, and an optional uncertainty head outputs standard deviations via the Softplus activation function.
@@
3.4. Selective State-Space (Mamba) Block
To efficiently capture long-range dependencies in the sequence, we employ a Selective State-Space Model (SSM) block, based on the Mamba architecture(Gu and Dao, 2023). An SSM is fundamentally a continuous-time linear system that models how a latent state $h(t)\in\mathbb{R}^{N}$ evolves based on an input signal $u(t)$, and how this state produces an output y(t)y(t)(Gu et al., 2022). The system is defined by:
@@ 3.5. Multi-Scale Temporal Convolutions
To capture local temporal patterns and contextual information at various resolutions, we incorporate a multi-scale temporal convolution module. This module operates on the hidden representations $H^{m}\in\mathbb{R}^{T\times d}$ from the preceding block. It employs a set of one-dimensional convolutional layers (Conv1D) with varying kernel sizes $k\in\mathcal{K}$ and dilation rates $\delta\in\mathcal{D}$. The kernel size $k$ determines the width of the local temporal window, while the dilation rate $\delta$ allows the receptive field to expand exponentially, enabling the capture of dependencies over longer ranges without a corresponding increase in computational cost.
For efficiency, we utilize depthwise separable convolutions, which factorize a standard convolution into a depthwise and a pointwise convolution(Wu et al., 2023). For each combination of kernel size $k$ and dilation $\delta$, a convolutional layer processes the input sequence:
@@
To dynamically integrate information from these different scales, we aggregate the resulting feature maps using an attention-style weighting mechanism(Vaswani et al., 2017). This allows the model to adaptively emphasize the most relevant temporal patterns for the given input. The aggregated representation $U^{m}$ is computed as a weighted sum:
@@ 3.6. Frequency-Domain Module
To capture global periodicities and trends, we employ a frequency-domain module that operates on the hidden sequence $H^{m}\in\mathbb{R}^{T\times d}$ from the previous block. This approach is inspired by classical signal processing and recent advances in time series forecasting that leverage frequency analysis(Zhou et al., 2022; Wu et al., 2021).
@@
Next, we perform spectral filtering by applying a learnable complex-valued spectral mask, $W_{f}^{m}\in\mathbb{C}^{T\times d}$, to the representation $Z^{m}$. This mask is parameterized either by its real and imaginary parts or by its magnitude and phase. The filtering operation is an element-wise (Hadamard) product:
@@
To enhance robustness and capture patterns across different time scales, we can optionally employ a multi-window analysis. Let $\mathcal{S}$ be a set of different analysis configurations (e.g., using different FFT window functions or segment lengths). For each configuration $s\in\mathcal{S}$, we compute a filtered time-domain representation $\widetilde{H}_{s}^{m}$. These representations are then aggregated using an adaptive weighting scheme:
@@
Finally, to ensure the stability and interpretability of the learned filter, we regularize the spectral mask $W_{f}^{m}$. The spectral gain is constrained by enforcing $\sup_{\omega}\lVert W_{f}^{m}(\omega)\rVert\leq G_{f}$, where $G_{f}$ is a predefined upper bound, implemented via parameterization and clipping. Additionally, we may apply penalties to encourage smoothness (e.g., penalizing the first-difference of the filter magnitudes) and sparsity (e.g., an L1 penalty on the magnitudes) to promote selective and stable spectral shaping.
@@ 3.7. Cross-Channel Attention
To explicitly model the interactions and dependencies between the $C$ different variables (channels) of the multivariate time series, we incorporate a cross-channel attention module. This allows the model to dynamically route information between related variables. The input to this module is a tensor $H\in\mathbb{R}^{T\times C\times d}$, representing the features for $C$ channels over $T$ time steps.
Following the standard self-attention mechanism(Vaswani et al., 2017), we first generate query, key, and value representations. However, to capture the relationships between channels rather than time steps, we compute attention scores based on a summary of each channel’s representation across time. Specifically, we first apply a temporal pooling operator, $\operatorname{pool}_{t}$, to the time series of each channel $H_{:,i,:}\in\mathbb{R}^{T\times d}$ to obtain a channel-specific summary vector $h_{i}^{\text{pool}}\in\mathbb{R}^{d}$. These summary vectors are then linearly projected to form the query ($q_{i}$) and key ($k_{j}$) vectors:
@@ 3.8. Adaptive Multi-Modal Ensemble
Our model processes the input signal through multiple, parallel feature extraction modules, including the Mamba-based model, multi-scale temporal convolutions, and the frequency-domain module. To synthesize insights from these diverse perspectives, we employ an adaptive ensemble that dynamically weights the contribution of each modality at every time step. This is crucial, as the predictive power of any single modality can fluctuate depending on the local characteristics of the time series.
First, each modality $m$ produces a feature representation, which we denote generically as $O^{m}\in\mathbb{R}^{T\times d}$. This representation is then passed to a dedicated prediction head, $f_{m}$, to generate a modality-specific forecast:
@@
Here, the colon (:) in the subscript of $O_{t,:}^{m}$ is a standard tensor notation indicating that we select all elements along the corresponding dimension. Specifically, $O_{t,:}^{m}$ extracts the complete feature vector of size $d$ for modality $m$ at time step $t$.
The core of the ensemble is the mechanism for generating the time-varying weights $w_{t}^{m}$, which functions as a form of attention mechanism. We first construct a context vector $c_{t}\in\mathbb{R}^{d_{c}}$ that summarizes the state of the system at time $t$ by aggregating information from all modalities. This is done by pooling the feature representations (e.g., via global average pooling), concatenating them, and processing the result through an MLP:
@@
Finally, to improve robustness under non-stationarity, we apply regularization to the weights $w_{t}^{m}$. Entropy regularization helps to mitigate overconfident collapse (where the model relies on a single modality), while temporal smoothness regularization penalizes erratic oscillations in the weights over time. This improves the stability and interpretability of the ensemble.

### v3

# Multi-modal Adaptive Estimation for Temporal Respiratory Disease Outbreak

@@ Abstract.
Timely and robust influenza incidence forecasting is critical for public health decision-making. This paper presents MAESTRO (Multi-modal Adaptive Estimation for Temporal Respiratory Disease Outbreak), a novel, unified framework that synergistically integrates advanced spectro-temporal modeling with multi-modal data fusion, including surveillance, web search trends, and meteorological data. By adaptively weighting heterogeneous data sources and decomposing complex time series patterns, the model achieves robust and accurate forecasts. Evaluated on over 11 years of Hong Kong influenza data (excluding the COVID-19 period), MAESTRO demonstrates state-of-the-art performance, achieving a superior model fit with an R² of 0.956. Extensive ablations confirm the significant contributions of its multi-modal and spectro-temporal components. The modular and reproducible pipeline is made publicly available to facilitate deployment and extension to other regions and pathogens, presenting a powerful tool for epidemiological forecasting.
@@
copyright: acmlicensedjournalyear: 2025doi: 10.48550/arXiv.2509.08578journal: TISTjournalvolume: 37journalnumber: 4article: 111publicationmonth: 8ccs: Applied computing Health informaticsccs: Computing methodologies Supervised learning by regressionccs: Computing methodologies Neural networksccs: Applied computing Forecastingccs: Mathematics of computing Time series analysis
@@ 1. Introduction
Seasonal influenza exerts a substantial burden on healthcare systems and society. Accurate short-horizon forecasting of influenza incidence is crucial for proactive resource allocation, timely public health alerts, and effective communication strategies. However, achieving such forecasts is challenging. Epidemiological data are often a complex blend of signals: traditional surveillance data can be noisy and lagged, while exogenous drivers like web search trends and meteorological conditions introduce distinct temporal dynamics and periodicities. This inherent heterogeneity demands a new class of models capable of not just processing multi-modal data, but also discerning and adapting to the most salient patterns across both time and frequency domains.
To address these challenges, this paper proposes MAESTRO (Multi-modal Adaptive Estimation for Temporal Respiratory Disease Outbreak). Rather than relying on a single monolithic architecture, MAESTRO is founded on a principled design that synergistically integrates specialized modules to tackle distinct aspects of the forecasting problem. It systematically decomposes time series to separate long-term trends from seasonal fluctuations and then processes these components through parallel pathways that include advanced state-space models for capturing long-range dependencies and frequency-domain analysis for identifying robust periodic patterns. This spectro-temporal approach allows the model to build a rich, multi-faceted understanding of the underlying dynamics. The insights from these diverse analytical perspectives are then dynamically synthesized through an adaptive ensemble mechanism, which intelligently weights each component’s contribution based on the evolving data context.
To validate the approach, MAESTRO was rigorously evaluated on a comprehensive real-world dataset from Hong Kong, spanning over a decade of surveillance, Google Trends, and meteorological data (excluding the COVID-19 pandemic period from training to ensure data integrity). Moreover, extensive ablation studies were conducted to systematically verify the contributions of each architectural component and data modality, providing a clear understanding of the model’s inner workings.
This work makes three primary contributions. First, a novel, unified spectro-temporal architecture is introduced that synergistically couples series decomposition, state-space modeling, and frequency-domain learning to specifically address the complexities of epidemiological forecasting. Second, an effective multi-modal integration strategy is proposed that cohesively merges heterogeneous data sources—epidemiological surveillance, web search trends, and weather—through cross-channel attention and an adaptive fusion framework. Third, through a comprehensive empirical study on influenza in Hong Kong, the model’s state-of-the-art accuracy and robustness are demonstrated, and full reproducibility is ensured by making the entire pipeline publicly available, thereby facilitating its adoption and extension to other diseases and regions.
@@ 2. Related Work
MAESTRO addresses this gap. It proposes a novel, modular architecture that synergistically combines time series decomposition, state-space modeling (Mamba), and frequency-domain learning. Crucially, it employs a cross-channel attention mechanism to robustly fuse multi-modal signals, positioning it as a comprehensive and powerful tool for modern epidemiological forecasting.
The principles of Multi-modal fusion and attention mechanisms, central to MAESTRO, have demonstrated significant value across a diverse range of applications beyond epidemiological forecasting. For instance, in industrial settings, methods combining LiDAR and camera data through multimodal fusion have been developed for precise defect detection (Li et al., 2025a). Similarly, attention mechanisms have been effectively utilized in complex detection tasks, such as identifying bird collision risks (Li and Yu, 2022) and enhancing object detection models for industrial quality control (Li et al., 2024). Furthermore, the fusion of advanced signal processing techniques like Adaptive Fourier Decomposition (AFD) with deep learning has shown promise in enhancing the interpretation of biomedical signals, such as spirometry data (Li et al., 2025b). These examples underscore the broad applicability and robustness of the core techniques that MAESTRO integrates and adapts for the specific challenges of forecasting respiratory diseases.
@@ 3.1. Problem Formulation
Given a multivariate time series $\mathbf{X}\in\mathbb{R}^{T\times D}$ (surveillance, trends, weather) and target $y\in\mathbb{R}$ (influenza positive rate), the goal is to predict $y_{t+1:t+H}$ from $\mathbf{X}_{t-L+1:t}$ with horizon $H$ and window $L$.
@@ 3.2. Notation and Assumptions
The elementwise (Hadamard) product is denoted by $\odot$ and convolution by $\ast$; $\mathcal{F}(\cdot)$ and $\mathcal{F}^{-1}(\cdot)$ are the 1D FFT and inverse FFT along time. Modalities are indexed by $m\in\{1,\dots,M\}$, channels by $i\in\{1,\dots,C\}$, and time by $t\in\{1,\dots,T\}$. Hidden width is $d$. The moving-average operator $\mathcal{D}_{K}$ is a length-$K$ low-pass filter (trend), and $I-\mathcal{D}_{K}$ yields the high-pass residual (seasonal). $\operatorname{softmax}$ is used for normalized exponentiation over the specified axis, $\operatorname{softplus}(x)=\log(1+e^{x})$ for nonnegativity, $\operatorname{pool}$ for temporal pooling (e.g., mean/max), and $\operatorname{MLP}$ for a small feed-forward network. Unless stated, $\|\cdot\|$ denotes the Euclidean/operator norm and $\rho(\cdot)$ the spectral radius.
@@ 3.3. Architecture Overview
    Series Decomposition: If the use_decomp flag is enabled, the input data is first decomposed into trend and seasonal components. This is achieved using an additive decomposition method based on a moving average filter, as implemented in the SeriesDecomposition module. Specifically, the trend component is extracted by applying a 1D average pooling layer with a specified kernel size. To handle boundary effects, the input sequence is padded by repeating the first and last values. The seasonal component is then obtained by subtracting the calculated trend from the original series.
@@
This modular and hierarchical architecture allows MAESTRO to effectively integrate diverse data modalities and capture a wide range of spectro-temporal dynamics. In summary, a moving-average block decomposes inputs into seasonal and trend components. These streams are then embedded and encoded with Transformer encoders. Residual Mamba state-space, multi-scale temporal convolutions, and a frequency-domain module (learnable filters in FFT space) are added. Cross-channel attention captures inter-variable dependencies. Finally, a linear temporal projection maps encoder outputs to the prediction horizon $H$, and an optional uncertainty head outputs standard deviations via the Softplus activation function.
@@
3.4. Mamba State-Space Model
To efficiently capture long-range dependencies in the sequence, a Mamba state-space model(Gu and Dao, 2023) is employed. An SSM is fundamentally a continuous-time linear system that models how a latent state $h(t)\in\mathbb{R}^{N}$ evolves based on an input signal $u(t)$, and how this state produces an output y(t)y(t)(Gu et al., 2022). The system is defined by:
@@ 3.5. Multi-Scale Temporal Convolutions
To capture local temporal patterns and contextual information at various resolutions, a multi-scale temporal convolution module is incorporated. This module operates on the hidden representations $H^{m}\in\mathbb{R}^{T\times d}$ from the preceding block. It employs a set of one-dimensional convolutional layers (Conv1D) with varying kernel sizes $k\in\mathcal{K}$ and dilation rates $\delta\in\mathcal{D}$. The kernel size $k$ determines the width of the local temporal window, while the dilation rate $\delta$ allows the receptive field to expand exponentially, enabling the capture of dependencies over longer ranges without a corresponding increase in computational cost.
For efficiency, depthwise separable convolutions are utilized, which factorize a standard convolution into a depthwise and a pointwise convolution(Wu et al., 2023). For each combination of kernel size $k$ and dilation $\delta$, a convolutional layer processes the input sequence:
@@
To dynamically integrate information from these different scales, the resulting feature maps are aggregated using an attention-style weighting mechanism(Vaswani et al., 2017). This allows the model to adaptively emphasize the most relevant temporal patterns for the given input. The aggregated representation $U^{m}$ is computed as a weighted sum:
@@ 3.6. Frequency-Domain Module
To capture global periodicities and trends, a frequency-domain module is employed that operates on the hidden sequence $H^{m}\in\mathbb{R}^{T\times d}$ from the previous block. This approach is inspired by classical signal processing and recent advances in time series forecasting that leverage frequency analysis(Zhou et al., 2022; Wu et al., 2021).
@@
Next, spectral filtering is performed by applying a learnable complex-valued spectral mask, $W_{f}^{m}\in\mathbb{C}^{T\times d}$, to the representation $Z^{m}$. This mask is parameterized either by its real and imaginary parts or by its magnitude and phase. The filtering operation is an element-wise (Hadamard) product:
@@
To enhance robustness and capture patterns across different time scales, a multi-window analysis can be optionally employed. Let $\mathcal{S}$ be a set of different analysis configurations (e.g., using different FFT window functions or segment lengths). For each configuration $s\in\mathcal{S}$, a filtered time-domain representation $\widetilde{H}_{s}^{m}$ is computed. These representations are then aggregated using an adaptive weighting scheme:
@@
Finally, to ensure the stability and interpretability of the learned filter, the spectral mask $W_{f}^{m}$ is regularized. The spectral gain is constrained by enforcing $\sup_{\omega}\lVert W_{f}^{m}(\omega)\rVert\leq G_{f}$, where $G_{f}$ is a predefined upper bound, implemented via parameterization and clipping. Additionally, penalties may be applied to encourage smoothness (e.g., penalizing the first-difference of the filter magnitudes) and sparsity (e.g., an L1 penalty on the magnitudes) to promote selective and stable spectral shaping.
@@ 3.7. Cross-Channel Attention
To explicitly model the interactions and dependencies between the $C$ different variables (channels) of the multivariate time series, a cross-channel attention module is incorporated. This allows the model to dynamically route information between related variables. The input to this module is a tensor $H\in\mathbb{R}^{T\times C\times d}$, representing the features for $C$ channels over $T$ time steps.
Following the standard self-attention mechanism(Vaswani et al., 2017), query, key, and value representations are first generated. However, to capture the relationships between channels rather than time steps, attention scores are computed based on a summary of each channel’s representation across time. Specifically, a temporal pooling operator, $\operatorname{pool}_{t}$, is first applied to the time series of each channel $H_{:,i,:}\in\mathbb{R}^{T\times d}$ to obtain a channel-specific summary vector $h_{i}^{\text{pool}}\in\mathbb{R}^{d}$. These summary vectors are then linearly projected to form the query ($q_{i}$) and key ($k_{j}$) vectors:
@@ 3.8. Adaptive Multi-modal Ensemble
The model processes the input signal through multiple, parallel feature extraction modules, including the Mamba-based model, multi-scale temporal convolutions, and the frequency-domain module. To synthesize insights from these diverse perspectives, an adaptive ensemble is employed that dynamically weights the contribution of each modality at every time step. This is crucial, as the predictive power of any single modality can fluctuate depending on the local characteristics of the time series.
First, each modality $m$ produces a feature representation, which is denoted generically as $O^{m}\in\mathbb{R}^{T\times d}$. This representation is then passed to a dedicated prediction head, $f_{m}$, to generate a modality-specific forecast:
@@
Here, the colon (:) in the subscript of $O_{t,:}^{m}$ is a standard tensor notation indicating that all elements along the corresponding dimension are selected. Specifically, $O_{t,:}^{m}$ extracts the complete feature vector of size $d$ for modality $m$ at time step $t$.
The core of the ensemble is the mechanism for generating the time-varying weights $w_{t}^{m}$, which functions as a form of attention mechanism. A context vector $c_{t}\in\mathbb{R}^{d_{c}}$ is first constructed that summarizes the state of the system at time $t$ by aggregating information from all modalities. This is done by pooling the feature representations (e.g., via global average pooling), concatenating them, and processing the result through an MLP:
@@
Finally, to improve robustness under non-stationarity, regularization is applied to the weights $w_{t}^{m}$. Entropy regularization helps to mitigate overconfident collapse (where the model relies on a single modality), while temporal smoothness regularization penalizes erratic oscillations in the weights over time. This improves the stability and interpretability of the ensemble.